OpenStudy (tiffany_rhodes):

initial value problem? Sin2xdx + cos6ydy = 0 , y(pi/2) = pi/6

OpenStudy (tiffany_rhodes):

I subtracted the cos6ydy to the other side and got sin2xdx = -cos6ydy. The I integrated both sides and got: -1/2cos2x +C = -1/6sin6y +C

OpenStudy (tiffany_rhodes):

I had trouble solving in terms of y

OpenStudy (tiffany_rhodes):

and then attempting to find c

OpenStudy (amoodarya):

\[ Sin2x dx + \cos6y dy = 0\rightarrow \cos 6y dy =- \sin 2x dx\] \[\int\limits \cos 6y dy =\int\limits- \sin 2x dx\] \[\frac{ 1 }{ 6 }\sin 6y =\frac{ 1 }{2 }\cos 2x +c\] now apply initial condition

OpenStudy (amoodarya):

(pi/6,pi/2) 1/6 sin (6 pi/2) =1/2 cos(2 pi/6 ) +c 0= 1/2 (1/2) +c so c=-1/4

OpenStudy (tiffany_rhodes):

Okay, so how would I express the equation in terms of y? So far I have sin(6y) = 3cos(2x) -1/4

OpenStudy (amoodarya):

it is not a function but if you restrict area may find y

OpenStudy (dumbcow):

in terms of y ... so you mean solve for x just use inverse cos \[\frac{1}{2}\cos 2x - \frac{1}{4} = \frac{1}{6} \sin 6y\] \[\cos 2x = \frac{1}{3} \sin 6y + \frac{1}{2}\] \[x = \frac{\cos^{-1} (\frac{1}{3} \sin 6y + \frac{1}{2})}{2}\] this will have restricted values for y as said by earlier post

OpenStudy (tiffany_rhodes):

I'm sorry, I meant solve for y in terms of x. So if you solved for y would the equation look like: \[y = (\sin^{-1} (3\cos2x -1/4))/(6) \] ?

OpenStudy (dumbcow):

yep wait distribute the 6 .... 3cos(2x) - 3/2 ...

OpenStudy (tiffany_rhodes):

so, \[y=\sin^{-1} (3\cos2x)-3/2\] ?

OpenStudy (dumbcow):

\[y = \frac{\sin^{-1} (3\cos 2x - \frac{3}{2})}{6}\]

OpenStudy (tiffany_rhodes):

ooh alright. thanks!

OpenStudy (amoodarya):

\[y=\sin^−1(3\cos2x−\frac{ 3 }{2 }) 6\rightarrow \frac{ -\pi }{ 2 } \le 3\cos2x−\frac{ 3 }{2 } \le \frac{ -\pi }{ 2 } \]

OpenStudy (amoodarya):

sorry right side is +pi/2